### v2

@@ Methods
Accuracy
We recommend a statistical comparison of the obtained spike data to reference data created by a trusted simulation code. The distributions of single-neuron firing rate (FR), distributions of coefficient of variation (CV) of the inter-spike intervals (ISI), and distributions of short-term spike-count correlation coefficients (CC) should be computed. For details, refer to [44, Sec. 2.3.1] and see also [96]. Demonstrating the visual overlap of these distributions with the reference is the minimum requirement. Computing scores to quantify the match of the distributions is appreciated such as Kullback-Leibler divergence, Kolmogorov-Smirnov test, or the Earth Mover Distance.
Performance
@@
Acknowledgements This project received funding from the European Union’s Horizon Europe Programme under the Specific Grant Agreement No. 101147319 (EBRAINS 2.0 Project); the Joint Lab “Supercomputing and Modeling for the Human Brain”; Juelich Research Centre intramural STEF fund for the update of instruments; HiRSE_PS, the Helmholtz Platform for Research Software Engineering - Preparatory Study, an innovation pool project of the Helmholtz Association; the Helmholtz Association’s Initiative and Networking Fund under project number SO-092 (Advanced Computing Architectures, ACA); the Federal Ministry of Education and Research (BMBF, Germany) through the projects NEUROTEC II (grant number 16ME0399) and Clusters4Future-NeuroSys (grant number 03ZU1106CA); EPSRC (grant numbers EP/P006094/1, EP/S030964/1 and EP/V052241/1); the European Union’s Horizon 2020 research and innovation program under Grant Agreement 945539 (HBP SGA3); and Google Summer of Code. The contribution of the University of Cagliari to this study was supported by the Project e.INS Ecosystem of Innovation for Next Generation Sardinia—spoke 10-CUP F53C22000430001—MUR Code No. ECS00000038. Authors involved in NEST GPU development acknowledge the use of Fenix Infrastructure resources, which are partially funded from the European Union’s Horizon 2020 research and innovation programme through the ICEI project under the Grant Agreement No. 800858.

### v3

@@ 4.2.2 Model and simulation parameters
4.2.3 Accuracy
We recommend a statistical comparison of the obtained spike data to reference data created by a trusted simulation code. The distributions of single-neuron firing rate (FR), distributions of coefficient of variation (CV) of the inter-spike intervals (ISI), and distributions of short-term spike-count correlation coefficients (CC) should be computed. For details, refer to [45, Sec. 2.3.1] and see also [98]. Demonstrating the visual overlap of these distributions with the reference is the minimum requirement. Computing scores to quantify the match of the distributions is appreciated such as Kullback-Leibler divergence, Kolmogorov-Smirnov test, or the Earth Mover Distance.
4.2.4 Performance
@@
This project received funding from the European Union’s Horizon Europe Programme under the Specific Grant Agreement No. 101147319 (EBRAINS 2.0 Project); the Joint Lab “Supercomputing and Modeling for the Human Brain”; Juelich Research Centre intramural STEF fund for the update of instruments; HiRSE_PS, the Helmholtz Platform for Research Software Engineering - Preparatory Study, an innovation pool project of the Helmholtz Association; the Helmholtz Association’s Initiative and Networking Fund under project number SO-092 (Advanced Computing Architectures, ACA); the Federal Ministry of Education and Research (BMBF, Germany) through the projects NEUROTEC II (grant number 16ME0399) and Clusters4Future-NeuroSys (grant number 03ZU1106CA); EPSRC (grant numbers EP/P006094/1, EP/S030964/1 and EP/V052241/1); the European Union’s Horizon 2020 research and innovation program under Grant Agreement 945539 (HBP SGA3); and Google Summer of Code. The contribution of the University of Cagliari to this study was supported by the Project e.INS Ecosystem of Innovation for Next Generation Sardinia—spoke 10-CUP F53C22000430001—MUR Code No. ECS00000038. Authors involved in NEST GPU development acknowledge the use of Fenix Infrastructure resources, which are partially funded from the European Union’s Horizon 2020 research and innovation programme through the ICEI project under the Grant Agreement No. 800858.